The Tops-For-All Store sells T-shirts: regular for $4, and deluxe for $5. Last month they sold 344 T-shirts. In addition, a third of the regular shirts and half of the deluxe shirts were imprinted, for $2 extra. Total gross reciepts were $1752. How many regular shirts were sold?
Answer by mananth(16946)

let regular shirts sold be x
deluxe shirt will be 344-x
..
regular shirt $4
deluxe shirt $5
.
sales from shirts = 4x+5(344-x)
..
imprint value
x/3 *2 + 1/2 * (344-x)
..
shirts sale + imprint value = total receipt
4x+5(344-x)+x/3 *2 + 1/2 * (344-x) = 1752
4x+1720 -5x +2x/3 +172-x/2= 1752
4x-5x+2x/3-x/2=1752-1720-172
-x+2x/3 -x/2 = -140
-6x+4x-3x / 6 = -140
-5x/6 = -140
-5x= 6*-140
x= 6*-140 / -5
x= 168 which is regular shirts
244-168= 76 deluxe shirts